OpenStudy (anonymous):

how Do you solve this .... What is the approximate volume of a cone with a height of 3 meters and base radius of 8 meters? Use 3.14

OpenStudy (anonymous):

\[1/3 \Pi r ^{2} h\]

OpenStudy (anonymous):

Ok that is what i thought , wasnt sure though thank you

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

solve it , please so i can see if i got it right ?

OpenStudy (anonymous):

why don't you tell me what you got first?

OpenStudy (anonymous):

i got 226.08 m3

OpenStudy (anonymous):

careful how you say to the third power. i think the better wording may be 200.96 meters cubed.

OpenStudy (anonymous):

true but is that right ?

OpenStudy (anonymous):

the reply you deleted had the correct number, just funny wording

OpenStudy (anonymous):

the current info has the correct power of three, but the wrong info.

OpenStudy (anonymous):

'but the wrong info' should say 'the wrong value'. sorry

OpenStudy (anonymous):

226.08 m cubed

OpenStudy (anonymous):

I am not seeing 226.08 as the correct value.

OpenStudy (anonymous):

its gotta b right

OpenStudy (anonymous):

show me your equation...

OpenStudy (anonymous):

8*3=24*3.14=75.36*3=226.08 m cubed

OpenStudy (anonymous):

hmm...\[1/3(\Pi)r ^{3}h\](1/3)x3.14x(8*8)x3

OpenStudy (anonymous):

i did exactly what u told me .

OpenStudy (anonymous):

where did you get the values for 8*3?

OpenStudy (anonymous):

let's take this one step at a time... what do you see as the radius?

OpenStudy (anonymous):

the radius was 8 & the height was 3 which equals 24 & then the pie (3.14) which is 75.36 & then cubed so Another three , which equaled 226.08 meter cubed (my mom said it looked right)

OpenStudy (anonymous):

where does the cubed come from?

OpenStudy (anonymous):

\[1/3 * \Pi * r ^{2} * h\]\[1/3 * 3.14 * 8^{2} * 3\] equals...\[1/3 * 3.14 * 64 * 3 = 200.96\]

OpenStudy (anonymous):

oK tHANK U